# FastVLM: Efficient Vision Encoding for Vision Language Models

### v1

@@ D.2 Visual Instruction Tuning Datasets
We use 3 different version of instruction tuning datasets. The smallest scale is LLaVA-1.5 665K dataset [49]. We further scale up this dataset by including training splits of the following datasets; AI2D [33], ScienceQA [55], ChartQA [57], COCO [46], DocVQA [59], DVQA [31], GeoQA+ [8], OCRVQA [61], SegmentAnything [35], SynthDoG-EN [34], TextVQA [68] and Visual Genome [36]. The conversational data for the listed datasets is sourced from [14]. The total number of samples in this dataset is 1.1 million and is referred to as “1.1M” in all the tables. We further scale-up instruction tuning dataset using image-based conversational data from Cambrian-7M [72], which amounts to 5.4 million samples. Filtered Cambrian-7M [72] is merged with “1.1M ” dataset to obtain “6.5M” instruction tuning dataset. We then append all available single-image instruction tuning data open-sourced by LLaVA-OneVision [40] to “6.5M” to obtain “11.9M” instruction tuning dataset. From Tab. 6, we see further improvements in VLM benchmarks when instruction tuning dataset is scaled, following trends exhibited by image encoders much bigger than FastViTHD.

### v2

@@ D.2 Visual Instruction Tuning Datasets
We use 3 different version of instruction tuning datasets. The smallest scale is LLaVA-1.5 665K dataset [53]. We further scale up this dataset by including training splits of the following datasets; AI2D [37], ScienceQA [59], ChartQA [63], COCO [51], DocVQA [65], DVQA [35], GeoQA+ [8], OCRVQA [67], SegmentAnything [39], SynthDoG-EN [38], TextVQA [74] and Visual Genome [40]. The conversational data for the listed datasets is sourced from [15]. The total number of samples in this dataset is 1.1 million and is referred to as “1.1M” in all the tables. We further scale-up instruction tuning dataset using image-based conversational data from Cambrian-7M [78], which amounts to 5.4 million samples. Filtered Cambrian-7M [78] is merged with “1.1M ” dataset to obtain “6.5M” instruction tuning dataset. We then append all available single-image instruction tuning data open-sourced by LLaVA-OneVision [45] to “6.5M” to obtain “11.9M” instruction tuning dataset. We then include roughly 0.6M samples from DocMatix [41] dataset to obtain “12.5M” instruction tuning dataset. From Tab. 6, we see further improvements in VLM benchmarks when instruction tuning dataset is scaled, following trends exhibited by image encoders much bigger than FastViTHD. The best performing models are achieved when Stage 2 models are further fine-tuned on the MammothVL instruction tuning dataset [30]. Specifically, we filter the corpus to include only single-image instruction tuning data, resulting in “10.6M” samples.
@@ D.3 Evaluations
E Qualitative Analysis
We analyzed failures across benchmarks and found: In text-rich benchmarks (e.g., DocVQA, ChartQA), failures occur when text is too small or precise alignment is needed (e.g., reading tables). In majority of cases where the text is too small, simply using higher input resolution can reduce errors as shown in Tab. 15. Some cases require broader general knowledge to obtain a correct response as seen in Tab. 14, in such cases using a bigger LLM reduces failures. Some cases require reasoning about a higher resolution image, in which case using a bigger LLM can decrease failures as seen in Tab. 16. Some failures result from misjudgment, where correct responses are misclassified by the LLM judge or incorrect labels, we ignore these cases in our analysis.
When to Use a Larger LLM
Dataset: MMMU [93]
User What is the common term for the yellow area surrounding the site of an infection? Options: [”I don’t know and I don’t want to guess”, ’Corona’, ’Border’, ’Halo’, ’Toxin zone’]
Ground Truth D
FastVLM-0.5B @ 256 E
FastVLM-0.5B @ 1024 E
FastVLM-1.5B @ 256 D
FastVLM-1.5B @ 1024 D
Dataset: MMMU [93]
User The sinoatrial (SA) node is indicated by ___. Options: [’A’, ’B’, ’C’, ’D’, ’E’]
Ground Truth A
FastVLM-0.5B @ 256 E
FastVLM-0.5B @ 1024 D
FastVLM-1.5B @ 256 A
FastVLM-1.5B @ 1024 A
Table 14: Cases that require broader general knowledge, in which case a larger LLM is preferred.
When to Use Higher Resolution
Dataset: GQA [34]
User What is sitting inside the bowls?
Ground Truth squash
FastVLM-0.5B @ 256 Sculpture
FastVLM-0.5B @ 1024 Squash
FastVLM-1.5B @ 256 Potato
FastVLM-1.5B @ 1024 Squash
Dataset: ChartQA [63]
User What was the highest expenditure on foreign military aid in 2009/10?
Ground Truth 3781
FastVLM-0.5B @ 256 Germany
FastVLM-0.5B @ 1024 3781
FastVLM-1.5B @ 256 275
FastVLM-1.5B @ 1024 3781
Table 15: Cases where increasing the resolution is sufficient to obtain a better response.
When to Use Higher Resolution and a Larger LLM
Dataset: ChartQA [63]
User What was the value of the commercial property market in 2016?
Ground Truth 883
FastVLM-0.5B @ 256 10000
FastVLM-0.5B @ 1024 871
FastVLM-1.5B @ 256 1100
FastVLM-1.5B @ 1024 883
Dataset: DocVQA [65]
User Under which department ‘Stockroom’ is organized?
Ground Truth Research Service Department
FastVLM-0.5B @ 256 Department of Chemistry
FastVLM-0.5B @ 1024 Library
FastVLM-1.5B @ 256 Research Department
FastVLM-1.5B @ 1024 Research Service Department
Table 16: Cases where increasing the resolution alone is not sufficient to obtain a better response. Larger LLM in combination with higher resolution is required to obtain a correct response.